OpenStudy (anonymous):

three times the difference of a number and seven is one less than twice the number find the number then solve

myininaya (myininaya):

3(n-7)=2n-1

OpenStudy (anonymous):

\[\color{red}{\text{myininaya will solve}}\]

OpenStudy (anonymous):

but you need to do step by step. that will help a lot

myininaya (myininaya):

3n-21=2n-1 3n-2n=-1+21 1n=20 n=20

OpenStudy (anonymous):

slowly my dear with explanation for each step will really help

myininaya (myininaya):

lol it explains itself

OpenStudy (anonymous):

if you say so.

myininaya (myininaya):

first i distribute the 3 on the left hand side second i subtracted 2n on both sides and added 21 on both sides lastly 1n=20 which is our answer since 1*n=n

OpenStudy (anonymous):

you see there is a lot of doing things to both sides meaning both sides of the equal sign to get what you want.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got the steps?

OpenStudy (anonymous):

\(3 \cdot\) (x - 7) = 2x - 1 three times difference of a number(x) and 7 is one less than twice the number So we have the equation: \[3(x-7) = 2x - 1\]

OpenStudy (anonymous):

That is the hardest part for most people to do.

myininaya (myininaya):

reading?

OpenStudy (anonymous):

Surprisingly yes. Reading comprehension is a challenge for a vast section of the population.

OpenStudy (anonymous):

It's getting better though.

myininaya (myininaya):

the only thing i can read is math lol

myininaya (myininaya):

= is always a replacement for is

myininaya (myininaya):

well unless you have is less than or is greater than

myininaya (myininaya):

is less than can be replaced with < is greater than can be replaced with > is (by itself) replaced with =